Weston Products manufactures an industrial cleaning compound that goes through three processing departments—Grinding, Mixing, and Cooking. All raw materials are introduced at the start of work in the Grinding Department. The Work in Process T-account for the Grinding Department for May is given below:

Work in Process—Grinding Department
Inventory, May 1 220,220 Completed and transferred
to the Mixing Department
?
Materials 525,610
Conversion 327,830
Inventory, May 31 ?

The May 1 work in process inventory consisted of 121,000 pounds with $131,890 in materials cost and $88,330 in conversion cost. The May 1 work in process inventory was 100% complete with respect to materials and 30% complete with respect to conversion. During May, 405,000 pounds were started into production. The May 31 inventory consisted of 91,000 pounds that were 100% complete with respect to materials and 60% complete with respect to conversion. The company uses the weighted-average method in its process costing system.

Required:

1. Compute the Grinding Department's equivalent units of production for materials and conversion in May.

2. Compute the Grinding Department's costs per equivalent unit for materials and conversion for May.

3. Compute the Grinding Department's cost of ending work in process inventory for materials, conversion, and in total for May.

4. Compute the Grinding Department's cost of units transferred out to the Mixing Department for materials, conversion, and in total for May.
Answer #1

Units completed and transferred= Work in process, May 1+Units started during May-Work in process, May 31

= 121000+405000-91000= 435000 pounds

1. Compute the Grinding Department's equivalent units of production for materials and conversion in May.

Materials Conversion costs
Units completed and transferred 435000 435000
Work in process, May 31 (91000*100%)= 91000 (91000*60%)= 54600
Equivalent units 526000 489600

2. Compute the Grinding Department's costs per equivalent unit for materials and conversion for May.

Materials Conversion costs
Cost of beginning work in process $131890 $88330
Cost added during May 525610 327830
Total costs $657500 $416160
Equivalent units 526000 489600
Cost per equivalent unit $1.25 $0.85

3. Compute the Grinding Department's cost of ending work in process inventory for materials, conversion, and in total for May.

Materials Conversion costs Total
Cost per equivalent units $1.25 $0.85
Ending equivalent units 91000 54600
Total cost of ending work in process $113750 $46410 $160160

4. Compute the Grinding Department's cost of units transferred out to the Mixing Department for materials, conversion, and in total for May.

Materials Conversion costs Total
Cost per equivalent units $1.25 $0.85
Units completed and transferred 435000 435000
Total cost of ending work in process $543750 $369750 $913500